OpenStudy (perl):

Find the exact coordinates of the centroid of the region bounded by y = 2x + 3, y = x^2

OpenStudy (shubhamsrg):

do you know the formulla directly ?

OpenStudy (shubhamsrg):

well leme tell then,, from the graph, 2x+3 will lie over x^2 in the bounded reigon.. now, let f(x) = 2x+3 and g(x) = x^2 let these 2 intersect at a and b (a < b) ( both are x-co-ordinates) let f(x) + g(x) =m and f(x) -g(x) =n also, integral (n dx) = A=area ,ofcorse ! now, there;s this direct formulla, which doesnt have a very tough derivation actually.. x co-ordinate is = integral (x*n dx) / A y cor-ordinate is =integral (n*m dx) /2A note: limits of all integrals above is from a to b. hope this helps..

OpenStudy (perl):

thanks , what next

OpenStudy (shubhamsrg):

substitute in the values..i guess??